write the equation of the circle in standard form. find the center, radius, intercepts, and graph the circle.
x^2+y^2-4x+18y=-69

Answer by stanbon(75887)

write the equation of the circle in standard form. find the center, radius, intercepts, and graph the circle.
x^2+y^2-4x+18y=-69
------
[x^2-4x+4] + [y^2+18y+81] = -69+4+81
------
(x-2)^2 + (y+9)^2 = 26
------
Center: (2,-9)
Radius: sqrt(26)
Intercepts:
x-int = ?
Let y = 0; solve for "x":
(x-2)^2 = 26-81 = -55
Therefore: no x-intercepts
----
y-int = ?
Let x = 0; solve for "y":
(y+9)^2 = 26-4
y+9 = +-sqrt(22)
y = -9+-sqrt(22)
=====================
The graph cannot be shown on this site.
Cheers,
Stan H.